# How regular is the evolute of a plane curve?

### v1

@@ 1. Introduction
In contrast to this, we are interested in the least possible amount of regularity to be required from the parent curve in order to get a certain regularity for the evolute, and conversely. At a first glance, it would seem that since the curvature depends on second derivatives and the evolute is expressed in terms of it, it should have two orders of regularity less than the parent curve. The result however is that the evolute has only one order of regularity less than the parent curve, provided that the curvature of that curve has non-vanishing derivative, see Theorem 1. This “bootstrap” phenomenon, an apparent regularity gain due to geometric constraints, reminds us of what happens when looking at distance functions and was pointed out in [5], who attribute the initial observation to the appendix of [4], see [7] for recent refinements.
@@
A plane curve is a continuous map $\gamma$ from an interval $(a;b)$ to the plane $\mathbb{R}^{2}$. For $k\in\mathbb{Z}_{+}$, $\alpha\in(0,1)$, we say the curve is of class $\mathcal{C}^{k}$ (resp. $\mathcal{C}^{k,\alpha}$) if its components are. In order to unify notation, we will write $\mathcal{C}^{k,0}$ for $\mathcal{C}^{k}$. We routinely use the fact that if $f,g\in\mathcal{C}^{k,\alpha}$, with $k\geq 1$, then $f\circ g\in\mathcal{C}^{k,\alpha}$ (this of course fails for $k=0,\alpha<1$).
@@
For a curve of class $\mathcal{C}^{1}$ and $t_{0}\in(a;b)$, the arc length is $s_{\gamma}(t):=\int_{t_{0}}^{t}|\gamma^{\prime}(u)|du$. Changing the point $t_{0}$ changes $s_{\gamma}(t)$ by an additive constant, and we will have to do that on occasion. When $\gamma$ is, in addition, regular, $s_{\gamma}$ is an increasing diffeomorphism from $(a;b)$ to $(s_{\gamma}(a),s_{\gamma}(b))$, of the same smoothness class as $\gamma$. We will consider $\gamma\circ s_{\gamma}^{-1}$ as the “same curve parametrized by arc length”, but will need to distinguish it in notation.
@@
When $k\geq 2$, $\kappa(s):=\varphi^{\prime}(s)$ is called the (signed) curvature of $\gamma$ and, if $\kappa(s)\neq 0$. Reducing the interval if needed, we make the standing assumption that for $t\in(a;b)$, $\kappa(t)\neq 0$. Then we can define $R(s):=1/\kappa(s)$ is the (signed) radius of curvature of $\gamma$. The well-known (plane) Frenet-Serret formulae immediately follow:
@@
    For $k\geq 3$, $0\leq\alpha\leq 1$: $\gamma$ is $\mathcal{C}^{k,\alpha}$-regular and its radius of curvature has a nonzero derivative, if and only if $\tilde{\gamma}$ is $\mathcal{C}^{k-1,\alpha}$-regular.
@@ 3. Local behavior
Let $\gamma:(a;b)\longrightarrow\mathbb{R}^{2}$ be a $\mathcal{C}^{2}$-regular curve such that its signed radius of curvature $R$ is strictly monotone increasing on $(a;b)$, denote its evolute by $\tilde{\gamma}$. Let $s_{1}\in(a;b)$.
@@
    Suppose that there is a strictly increasing function $\omega:[0,\infty)\longrightarrow[0,\infty)$ with $\omega(0)=0$ such that for $s$ in a neighborhood of $s_{1}$, $|R(s)-R(s_{1})|\geq\omega(|s-s_{1}|)$. Then $\tilde{\gamma}$ can be represented in the local frame by $y=g(x)$, where $g(x)=O(x\omega^{-1}(x))$.
@@
    In particular, if $\gamma$ is $\mathcal{C}^{k}$-regular and $1<m:=m_{R}(s_{1})\leq k$, $\tilde{\gamma}$ is $\mathcal{C}^{1,1/m}$-regular near $s_{1}$ with $m\geq 3$, and cannot be any smoother.
@@
(5) $\langle\tilde{\gamma}(s_{2})-\tilde{\gamma}(s_{1}),\nu(s_{1})\rangle=(R(s_{2})-R(s_{1}))+\int_{s_{1}}^{s_{2}}\frac{R(s)-R(s_{2})}{R(s)}\langle\gamma^{\prime}(s)-\gamma^{\prime}(s_{1}),\nu(s_{1})\rangle ds\\ =(R(s_{2})-R(s_{1}))(1+O(|s_{2}-s_{1}|)),\\ \langle\tilde{\gamma}(s_{2})-\tilde{\gamma}(s_{1}),\gamma^{\prime}(s_{1})\rangle=\int_{s_{1}}^{s_{2}}\frac{R(s)-R(s_{2})}{R(s)}\langle\gamma^{\prime}(s),\gamma^{\prime}(s_{1})\rangle ds\\ =O(|R(s_{2})-R(s_{1})||s_{2}-s_{1}|).$
This implies in particular that the projection from the image of $\tilde{\gamma}$ to the $x$-axis is one-to-one, therefore we can represent that image curve as a graph $y=g(x)$; and since $x\asymp R(s_{2})-R(s_{1})\geq\omega(s_{2}-s_{1})$ and $|y|\preceq|R(s_{2})-R(s_{1})||s_{2}-s_{1}|$, we have $|y|\preceq x\omega^{-1}(x)$.
@@
Since $\gamma$, parametrized by arclength, is (at least) $\mathcal{C}^{2}$-regular, we can write $s=\gamma^{-1}(\gamma(s))$, where $\gamma^{-1}$ is of class $\mathcal{C}^{2}$ and only defined locally on the embedded manifold $\gamma(I)$, $I$ a small enough interval. However $\gamma$ is also given as an involute of $\tilde{\gamma}$ by $\gamma_{1}(\tilde{s})=\tilde{\gamma}_{1}(\tilde{s})-\tilde{\gamma}_{1}^{\prime}(\tilde{s})(s-c)$, where $\tilde{s}$ is arclength on the evolute $\tilde{\gamma}$. By the regularity hypothesis, $\gamma_{1}$ will be of class $\mathcal{C}^{0,\alpha}$ and $)s=\gamma^{-1}(\gamma_{1}(\tilde{s}$)), so that there is a constant $C>0$ such that
$|s_{2}-s_{1}|\leq C|\tilde{s}_{2}-\tilde{s}_{1}|^{\alpha}=C|R(s_{2})-R(s_{1})|^{\alpha}\leq C_{\varepsilon}^{\prime}|s_{2}-s_{1}|^{\alpha(m_{R}(s_{1})^{-}-\varepsilon)}$
for any $\varepsilon>0$, so $\alpha m_{R}(s_{1})^{-}\leq 1$.
@@
The hypothesis implies that $R^{\prime}$ admits an isolated zero at $s_{1}$, so on a punctured neighborhood of $s_{1}$, $\tilde{\gamma}$ is of class at least $\mathcal{C}^{2}$ since $k\geq 3$. We can take $\omega(x)=Cx^{m}$ and deduce from (i) that $\tilde{\gamma}$ can be represented in the local frame by $y=g(x)$ et $g(x)=O(|x|^{1+1/m})$, in particular $\tilde{\gamma}$ is differentiable at $s_{1}$ with $\tilde{\gamma}_{1}^{\prime}(\tilde{s}_{1})=\nu(s_{1})$. More generally, $\tilde{\gamma}^{\prime}(R(s))=\nu(s)$ in a neighborhood of $s_{1}$ and
$|\nu(s_{2})-\nu(s_{1})|\preceq|s_{2}-s_{1}|\asymp|R(s_{2})-R(s_{1})|^{1/m},$
by the hypothesis on the order of $R$ at $s_{1}$.
@@
We already know by the proof of Theorem 1 (i) that $R$ must be strictly monotone. If $R^{\prime}$ vanished at $s_{1}$, we would have to have $m_{R}(s_{1})\geq 3$, so $\tilde{\gamma}$ is at best of class $\mathcal{C}^{1,1/3}$ by Proposition 4(iii) and cannot be of class $\mathcal{C}^{2}$. $\Box$

### v2

@@ 1. Introduction
In contrast to this, we are interested in the least possible amount of smoothness to be required from the parent curve in order to get a certain smoothness for the evolute, and conversely. At a first glance, it would seem that since the curvature depends on second derivatives and the evolute is expressed in terms of it, it should have two orders of regularity less than the parent curve. The result however is that the evolute has only one order of regularity less than the parent curve, provided that the curvature of that curve has non-vanishing derivative, see Theorem 1. This “bootstrap” phenomenon, an apparent regularity gain due to geometric constraints, reminds us of what happens when looking at distance functions and was pointed out in [5], who attribute the initial observation to the appendix of [4], see [7] for recent refinements.
@@
A plane curve is a continuous map $\gamma$ from an interval $(a;b)$ to the plane $\mathbb{R}^{2}$. For $k\in\mathbb{Z}_{+}$, $\alpha\in(0,1)$, we say the curve is of class $\mathcal{C}^{k}$ (resp. $\mathcal{C}^{k,\alpha}$) if its components are. In order to unify notation, we will write $\mathcal{C}^{k,0}$ for $\mathcal{C}^{k}$. Recall that a function is of class $\mathcal{C}^{\infty}$ if it is of class $\mathcal{C}^{k}$ for any $k$. We routinely use the fact that if $f,g\in\mathcal{C}^{k,\alpha}$, with $k\geq 1$, then $f\circ g\in\mathcal{C}^{k,\alpha}$ (this of course fails for $k=0,\alpha<1$).
@@
For a curve of class $\mathcal{C}^{1}$ and $t_{0}\in(a;b)$, the arc length is $s_{\gamma}(t):=\int_{t_{0}}^{t}|\gamma^{\prime}(u)|du$. Changing the point $t_{0}$ changes $s_{\gamma}(t)$ by an additive constant, and we will have to do that on occasion. When $\gamma$ is, in addition, regular, $s_{\gamma}$ is an increasing diffeomorphism from $(a;b)$ to $(s_{\gamma}(a),s_{\gamma}(b))$, of at least the same smoothness class as $\gamma$; the arclength parametrization therefore provides the best possible smoothness for the curve among the regular parametrizations. We will consider $\gamma\circ s_{\gamma}^{-1}$ as the “same curve parametrized by arc length”, but will need to distinguish it in notation.
@@
When $k\geq 2$, $\kappa(s):=\varphi^{\prime}(s)$ is called the (signed) curvature of $\gamma$. Reducing the interval if needed, we make the standing assumption that for $s\in(a;b)$, $\kappa(s)\neq 0$. Then we can define $R(s):=1/\kappa(s)$ is the (signed) radius of curvature of $\gamma$. The well-known (plane) Frenet-Serret formulae immediately follow:
@@
    For $k\geq 3$, $0\leq\alpha\leq 1$: $\gamma$ is $\mathcal{C}^{k,\alpha}$-regular and $R^{\prime}(s)\neq 0$, if and only if $\tilde{\gamma}$ is $\mathcal{C}^{k-1,\alpha}$-regular.
@@ 3. Local behavior
Let $\gamma:(a;b)\longrightarrow\mathbb{R}^{2}$ be a $\mathcal{C}^{2}$-regular curve, denote its evolute by $\tilde{\gamma}$. Let $s_{1}\in(a;b)$.
@@
    Suppose that the signed radius of curvature $R$ of $\gamma$ is strictly increasing on $(a;b)$ and let $\omega(x):=R(s_{1}+x)-R(s_{1})$. Then $\tilde{\gamma}$ can be represented in the local frame by $y=g(x)$, where $g(x)=O(x\omega^{-1}(x))$.
@@
    If the derivative of $R$ vanishes to infinite order at $s_{1}$, the curve $\gamma$ is of class $\mathcal{C}^{1,0}$, and no better in the $\mathcal{C}^{1,\alpha}$ scale.
@@
    In particular, if $\gamma$ is $\mathcal{C}^{k}$-regular, if $R$ is strictly increasing on $(a;b)$ and $1<m:=m_{R}(s_{1})\leq k$, $\tilde{\gamma}$ is $\mathcal{C}^{1,1/m}$-regular near $s_{1}$ with $m\geq 3$, and cannot be any smoother.
@@
Notice that in item (i) we could have used any strictly increasing $\omega$ such that for $s$ in a neighborhood of $s_{1}$, $|R(s)-R(s_{1})|\geq\omega(|s-s_{1}|)$. This can apply to cases where $R$ is not monotone.
@@
(5) $\langle\tilde{\gamma}(s_{2})-\tilde{\gamma}(s_{1}),\nu(s_{1})\rangle=(R(s_{2})-R(s_{1}))+\int_{s_{1}}^{s_{2}}\frac{R(s)-R(s_{2})}{R(s)}\langle\gamma^{\prime}(s)-\gamma^{\prime}(s_{1}),\nu(s_{1})\rangle ds\\ =(R(s_{2})-R(s_{1}))(1+O(|s_{2}-s_{1}|^{2})),\\ \langle\tilde{\gamma}(s_{2})-\tilde{\gamma}(s_{1}),\gamma^{\prime}(s_{1})\rangle=\int_{s_{1}}^{s_{2}}\frac{R(s)-R(s_{2})}{R(s)}\langle\gamma^{\prime}(s),\gamma^{\prime}(s_{1})\rangle ds\\ =O(|R(s_{2})-R(s_{1})||s_{2}-s_{1}|).$
This implies in particular that the projection from the image of $\tilde{\gamma}$ to the $x$-axis is one-to-one, therefore we can represent that image curve as a graph $y=g(x)$; and since $x\asymp R(s_{2})-R(s_{1})$, $R(s_{2})-R(s_{1})\geq\omega(s_{2}-s_{1})$ for $s_{2}>s_{1}$, $R(s_{2})-R(s_{1})\leq-\omega(s_{2}-s_{1})$ for $s_{2}<s_{1}$, and $|y|\preceq|R(s_{2})-R(s_{1})||s_{2}-s_{1}|$, we have $|y|\preceq x\omega^{-1}(x)$.
@@
Since $\gamma$, parametrized by arclength, is (at least) $\mathcal{C}^{2}$-regular, we can write $s=\gamma^{-1}(\gamma(s))$, where $\gamma^{-1}$ is of class $\mathcal{C}^{2}$ and only defined locally on the embedded manifold $\gamma(I)$, $I$ a small enough interval. However $\gamma$ is also given as an involute of $\tilde{\gamma}$ by $\gamma_{1}(\tilde{s})=\tilde{\gamma}_{1}(\tilde{s})-\tilde{\gamma}_{1}^{\prime}(\tilde{s})(s-c)$, where $\tilde{s}$ is arclength on the evolute $\tilde{\gamma}$. By the regularity hypothesis, $\gamma_{1}$ will be of class $\mathcal{C}^{0,\alpha}$ and $)s=\gamma^{-1}(\gamma_{1}(\tilde{s}$)), so there is a constant $C>0$ such that
$|s_{2}-s_{1}|\leq C|\tilde{s}_{2}-\tilde{s}_{1}|^{\alpha}=C|R(s_{2})-R(s_{1})|^{\alpha};$
if $m_{R}(s_{1})^{-}<\infty$, for any $\varepsilon>0$ we have $C_{\varepsilon}$ such that $|R(s_{2})-R(s_{1})|\leq C_{\varepsilon}|s_{2}-s_{1}|^{m_{R}(s_{1})^{-}-\varepsilon}$, so that $|s_{2}-s_{1}|\leq C_{\varepsilon}^{\prime}|s_{2}-s_{1}|^{\alpha(m_{R}(s_{1})^{-}-\varepsilon)}$, so $\alpha m_{R}(s_{1})^{-}\leq 1$.
If $m_{R}(s_{1})^{-}=\infty$, $|R(s_{2})-R(s_{1})|\leq C|s_{2}-s_{1}|^{1+1/\alpha}$, which leads to a contradiction.
For the second statement, Theorem 1 implies that $R$ is of class $\mathcal{C}^{1}$. If $R$ was of class $\mathcal{C}^{1,\alpha}$, $\alpha>0$, we would have $m_{R}(s_{1})^{-}\leq 1/\alpha$ which contradicts the hypothesis of vanishing of the derivative to infinite order.
@@
The hypothesis implies that $R^{\prime}$ admits an isolated zero at $s_{1}$, so on a punctured neighborhood of $s_{1}$, $\tilde{\gamma}$ is of class at least $\mathcal{C}^{2}$ since $k\geq 3$. We have $\omega(x)\asymp x^{m}$ and deduce from (i) that $\tilde{\gamma}$ can be represented in the local frame by $y=g(x)$ et $g(x)=O(|x|^{1+1/m})$, in particular $\tilde{\gamma}$ is differentiable at $s_{1}$ with $\tilde{\gamma}_{1}^{\prime}(\tilde{s}_{1})=\nu(s_{1})$.
More generally, for $s,s^{\prime}$ in a neighborhood of $s_{1}$, $\tilde{\gamma}^{\prime}(R(s))=\nu(s)$ and $|\nu(s^{\prime})-\nu(s_{1})|\preceq|s^{\prime}-s|$; we claim that
$|s^{\prime}-s|\preceq|R(s^{\prime})-R(s)|^{1/m}.$
To see this, assume without loss of generality that $R^{(m)}(t)\geq c>0$ for $t$ in an open interval $I$ around $s_{1}$, and $s,s^{\prime}\in I$, $s_{1}\leq s<s^{\prime}$. An easy induction argument then shows that for $0\leq j\leq m-1$, $R^{(m-j)}(s)\geq\frac{c}{j!}(s-s_{1})^{j}\geq 0$. By the Taylor-Lagrange formula with integral remainder,
$R(s^{\prime})-R(s)=\sum_{j=1}^{m-1}\frac{(s^{\prime}-s)^{j}}{j!}R^{(j)}(s)+\int_{s}^{s^{\prime}}\frac{(s^{\prime}-t)^{m-1}}{(m-1)!}R^{(m)}(t)dt\\ \geq c\int_{s}^{s^{\prime}}\frac{(s^{\prime}-t)^{m-1}}{(m-1)!}dt=c\frac{(s^{\prime}-s)^{m}}{m!}.$
This proves the claim.
@@
We already know by the proof of Theorem 1 (i) that $R$ must be strictly monotone. If $R^{\prime}$ vanished at $s_{1}$, we would have to have $m_{R}(s_{1})\geq 3$, so $\tilde{\gamma}$ is at best of class $\mathcal{C}^{1,1/3}$ by Proposition 4(ii) and (iii), and cannot be of class $\mathcal{C}^{2}$. $\Box$